On October 17, 1973, the Organization of Arab Petroleum Exporting Countries (OPEC) declared an embargo on the
shipment of oil to those countries that had supported Israel in its conflict with Egypt. With one stroke, the total
dependence of the industrialized world on oil, especially the US, became painfully clear. The effects of the embargo
were immediate. The retail price of a gallon of gasoline rose from a national average of 40 cents in May 1973 to 60
cents in June 1974. President Nixon (a controlling man that he was) decided to set a price ceiling at 50 cents.
Illustrate using supply and demand graph the effect on equilibrium quantity and quantity sold of gasoline as the
embargo came into effect. And then after the embargo, (on the same graph) show what happens to the quantity of
gasoline sold as Nixon imposes the price ceiling. Is there any dead weight loss from the above two events?
1
Expert's answer
2020-04-30T09:52:56-0400


Before embargo there was an equlibrium - (40;Q*).

When embardo is imposed the equilibrium quantity decreases to Q' and the price rises to 60.

After price ceiling establishment we see on the graf that the price decreases to 50, but the gap between Qs and Qd, which are quantity supplied and quantity demanded, occurs. So, at price of 50 cents the quantity sold/purchased will be Qs, which is less that before price ceiling and before embargo. Thus, we see the decrease in welfare.

After embargo:

Consumer surplus = a+b.

Supplier surplus = c+d+e.

After price ceilling:

Cons.surplus = a+b.

Suppl.surplus= c.

So, the dead weight loss is (d+e).